### v2

# EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes

Abstract
For decades, classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over such classical models. However, the currently-used benchmark for NPPs does not represent an up-to-date challenge in the seismological community, since it contains data leakage and omits the largest earthquake sequence from the region. Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology. To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by an implementation of the state-of-the-art forecasting model: ETAS. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. Benchmarking experiments, using both log-likelihood and generative evaluation metrics widely recognised in seismology, show that none of the five NPPs tested outperform ETAS. These findings suggest that current NPP implementations are not yet suitable for practical earthquake forecasting. Nonetheless, EarthquakeNPP provides a platform to foster future collaboration between the seismology and machine learning communities.
1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model [53, 54], despite significant growth in available data [79, 72, 62, 90, 51, 80, 50].
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility [9, 56, 69, 31, 5, 97, 96]. While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model. These datasets are derived from publicly available raw data, which we process and configure within our platform to facilitate meaningful forecasting experiments relevant to stakeholders in the seismology community. Covering various regions of California, these datasets represent typical forecasting zones and encompass data commonly utilized by forecast issuers. Moreover, employing modern techniques, some datasets include smaller magnitude earthquakes, exploring the potential of numerous small events to enhance forecasting performance through flexible NPPs. To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as the benchmark for NPPs.
Although initial benchmarking shows that none of the five tested NPP implementations outperform ETAS, EarthquakeNPP is designed to support ongoing model development and evaluation. In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking. This ensures that future NPPs (and other models such as time series approaches [88] and Bayesian point processes [68]) can have direct relevance to seismological stakeholders. All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP/tree/main.
1.1 Related Work
Benchmarking by the NPP Community. Chen et al. [5] introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake [49], the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
The dataset segments are divided for training, testing, and validation. Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern. This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering [13] - a form of data leakage. Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards in time.
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. [5], following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
Chen et al. [5] benchmark their model against another spatio-temporal model, Neural Jump SDEs [31], and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark [97, 94, 96] similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
Benchmarking by the Seismology Community. Model comparison has been crucial in the development of earthquake forecasting models since their inception [34, 53]. The Collaboratory for the Study of Earthquake Predictability (CSEP) [41, 66, 64, 30] [https://cseptesting.org/ ] aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. [8] extends a temporal-only NPP from Shchur et al. [69] to include earthquake magnitudes. The second by Stockman et al. [77] extends another temporal-only model by Omi et al. [56] to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
2 Background
2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{\Delta t|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
where $\mathcal{H}_{t}=\{(t_{i},\mathbf{x}_{i})|t_{i}<t\}$ denotes the history of events preceding time $t$ and $|B(\mathbf{x},\Delta\mathbf{x})|$ is the Lebesgue measure of the ball $B(\mathbf{x},\Delta\mathbf{x})$ with radius $\Delta\mathbf{x}$. Given we observe a history of events up to $t_{i}$, the probability density function (pdf) of observing an event at time $t$ and location $\mathbf{x}$ is given by
$p(t,\mathbf{x}|\mathcal{H}_{t_{i}})=\lambda(t,\mathbf{x}|\mathcal{H}_{t_{i}})\cdot\exp\left(-\int_{t_{i}}^{t}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds\right).$ (2)
Most models specify the conditional intensity function, though some [e.g. 69, 5, 94] directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
$\log p(\mathcal{H}_{T})=\underbrace{\sum_{i=0}^{n}\log\lambda(t_{i}|\mathcal{H}_{t_{i}})-\int_{T_{0}}^{T_{1}}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds}_{\text{Temporal log-likelihood}}+\underbrace{\sum_{i=0}^{n}\log f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})}_{\text{Spatial log-likelihood}},$ (3)
where the decomposition of the spatio-temporal conditional intensity function, $\lambda(t_{i},\mathbf{x}_{i}|\mathcal{H}_{t_{i}})=\lambda(t_{i}|\mathcal{H}_{t_{i}})\cdot f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})$, allows the log-likelihood to be written as contributions from the temporal and spatial components. In practice, this exact function is often not maximized directly during training: for models specified through the conditional intensity function, an analytical solution to the integral term is generally not possible and is approximated numerically.
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community [see 95, and references therein] as well as the wider general point process literature [7], which now includes neural point processes [70]. The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the difference in mean log-likelihood per event provides the information gain from one model to another [7].
Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting [47] and established benchmarking experiments (CSEP) [81, 60]. The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time [33]. Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin — a delay that could miss critical, potentially damaging events.
2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model [54] is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California [44], New-Zealand [6], Italy [74], Japan [57] and Switzerland [48], and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments [e.g. 93, 60, 81, 4, 38, 39, 40]. The general formulation of the model is
[…]
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering [3] and static stress triggering [16, 39].
Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$. Veen and Schoenberg [87] developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
$\mathbf{\theta}^{(k+1)}=\operatorname*{arg\,max}_{\mathbf{\theta}}\mathbb{E}_{\mathbf{B}\sim p(\cdot|\mathcal{H}_{T_{1}},\mathbf{\theta}^{(k)})}\left[\log p(\mathcal{H}_{T_{1}},\mathbf{B}|\mathbf{\theta})\right].$ (5)
This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation, and simultaneously distinguishes background events from triggered events.
Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP. Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans. In red is a fault map from the GEM Global Active Faults Database [78].
The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. [46]. It defines the triggering kernel as
$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{c})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{c})}\right)^{1+\rho}},$ (6)
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$. This triggering kernel is derived from statistical distributions found through decades of observational studies [84, 83, 85] and several of the learnable parameters have been linked to physical properties of the earthquake rupture process [85, 29].
3 EarthquakeNPP Datasets
The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms [14, 11, 12]. It encompasses the San Andreas fault plate boundary system [99] and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://github.com/ss15859/EarthquakeNPP/tree/main, accompanied by more detailed documentation for each dataset. A summary of how earthquake datasets are generated, along with the associated challenges of using earthquake catalog data can be found in Appendix A. Table 1 provides a short summary of each EarthquakeNPP dataset.
Table 1: Summary of EarthquakeNPP datasets, including: region, dataset development, magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training (combined with validation) events, and number of testing events. The chronological partitioning of training, validation, and testing periods is also detailed. An auxiliary (burn-in) period begins from the Start date, followed by the respective starts of the training, validation, and testing periods. All dates are given as 00:$0$ UTC on January $1^{\text{st}}$, unless noted (* refers to 00:$0$ UTC on January $17^{\text{th}}$). Finally, we give our purpose for including each dataset.
ComCat SCEDC White QTM
Region Whole of California Southern California San Jacinto Fault-Zone QTM_SanJac:
San Jacinto Fault-Zone,
QTM_SaltonSea: Salton Sea
Development The U.S. Geological Survey (USGS) National Earthquake Information Center (NEIC) monitors global earthquakes (Mw 4.5 or larger) and provides complete seismic monitoring of the United States for all significant earthquakes (> Mw 3.0 or felt). Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products. The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California [28] since the Caltech Seismological Laboratory began routine operations in 1932. Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8. White et al. [90] created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California. This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion [90]. Using data collected by the SCSN, Ross et al. [62] generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
$\mathbf{M}_{\mathbf{c}}$ Mw 2.5 SCEDC_20: Mw 2.0,
SCEDC_25: Mw 2.5,
SCEDC_30: Mw 3.0
Mw 0.6 Mw 1.0
# Train/Test
Events
79,037 / 23,059 SCEDC_20:
128,265 / 14,351,
SCEDC_25:
43,221 / 5,466,
SCEDC_30:
12,426 / 2,065
38,556 / 26,914 QTM_SanJac:
18,664 / 4,837,
QTM_SanJac:
44,042 / 4,393
Start-Train-Val-Test-End 1971-1981-1998-2007-2020 1981-1985-2005-2014-2020 2008-2009-2014-2017-2021 2008-2009-2014-2016-2018
Purpose Example of data currently in use for operational
forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public.)
Three magnitude thresholds (Mw 2.0, 2.5, 3.0) explore the effect of truncation on forecasting model performance. To explore if newly detected low magnitude earthquakes contain additional predictive information. To explore if newly detected low magnitude earthquakes contain additional predictive information (with different detection methodology).
4 Benchmarking Experiment
We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs), each with prior positive claims in earthquake forecasting, against the ETAS model described in Section 2.2.
Neural Spatio-Temporal Point Process (NSTPP) [5]: A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs). We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model [5].
Deep Spatio-Temporal Point Process (DeepSTPP) [97]: A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process. This model features a closed-form intensity function, eliminating the need for numerical approximations.
Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) [96]: A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
Spatio-temporal Diffusion Point Process (DSTPP) [94]: A generative NPP that does not have a likelihood function. DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)[36]: A generative NPP that also lacks a likelihood function. SMASH adopts a normalization-free objective by estimating the pseudolikelihood of marked STPPs through score-matching.
4.1 Likelihood Evaluation
Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development [53, 23]. Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting [27], as earthquake occurrence follows power law distributions [33, 10] that are heavy-tailed, making the errors non-Gaussian and non-Laplacian — contrary to the assumptions underlying RMSE and MAE (see Section G).
For the three models with valid likelihood functions (NSTPP, DeepSTPP, and AutoSTPP), we present the mean log-likelihood scores in Figures 2 and 3. These scores are compared alongside the ETAS model (Section 2.2) and a homogeneous Poisson process. The Poisson model is fit to events in the auxiliary, training, and validation windows to provide a baseline score for comparison.
Unlike the NPPs, which follow the standard machine learning procedure of training, validation, and testing, ETAS does not typically incorporate validation in its estimation procedure. Thus, it is fit using both the training and validation windows combined. For NPPs, the likelihood for training, validation, and testing depends on events occurring before the respective splits through memory in their history. The exception is NSTPP, which lacks a direct dependency on prior events. Nevertheless, its likelihood is evaluated on the same data as the other models. The definition of the ETAS model (equation 4) specifies how the magnitudes of earthquakes in the history contribute towards the intensity function. This earthquake magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modeling choices beyond the scope of this work.
The ETAS model consistently achieves the highest temporal log-likelihood, with NPPs performing comparably or, in some cases, marginally better, except at the higher magnitude thresholds of the SCEDC catalog. Among the NPPs, AutoSTPP and NSTPP perform well across several datasets, though their performance is more variable than that of DeepSTPP, which demonstrates consistent, comparable performance to ETAS. Closer examination of model performance over time (Figure 9) reveals that relative performance to ETAS is poorest during large earthquake sequences. This is likely due to ETAS leveraging the magnitude feature of the data, which enables it to handle these sequences effectively. Conversely, model performance is strongest during "background" periods, when no large earthquakes occur. During these periods, ETAS models the background with a constant rate, while the NPPs improve upon this by capturing the non-stationary nature of the background data. This effect is most pronounced in the ComCat dataset, which includes more complex physical processes, such as those near the Mendocino Triple Junction [24]. The improved relative temporal performance of all NPPs compared to ETAS, particularly when the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
ETAS consistently outperforms all NPPs in spatial log-likelihood. As with temporal performance, relative performance to ETAS is weakest during large earthquake periods, likely due to the absence of a magnitude feature in the NPPs. AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu [96]), which are commonly observed in earthquake data [58].
Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
4.2 CSEP Consistency Tests
EarthquakeNPP supports the earthquake forecast evaluation protocol developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). In this procedure, a model generates 24-hour forecasts through 10,000 repeated simulations of earthquake sequences at the beginning of each day in the testing period. This approach mirrors how earthquake forecasts are produced in operational settings [86]. Models are then evaluated by comparing the observed sequence with the distribution of forecasts generated by the simulations. Four test statistics assess the temporal, spatial, likelihood, and magnitude components of the forecasts. A test is considered failed if the observed statistic falls within a pre-defined rejection region (Figure 11). We apply this evaluation procedure to the two generative NPPs (DSTPP and SMASH) alongside ETAS (Table 2) and present a case study on the 2010 M7.2 El Mayor-Cucapah earthquake, using the forecasts from these models (Figure 4). Appendix E provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/.
Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets. A test is performed at the $\alpha=0.05$ significance level on each day in the testing period. The pass rate indicates the proportion of testing days with non-rejected hypotheses. If the model is the data generator, quantile scores should be uniformly distributed. The KS-Statistic quantifies deviation from uniformity (see Appendix E). ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
Dataset Model Number Test Spatial Test Pseudo Likelihood Test Magnitude Test
Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat.
ComCat ETAS 85.1% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
SMASH 72.4% 0.212 68.6% 0.217 87.6% 0.134
DSTPP 70.4% 0.116 88.6% 0.070 86.3% 0.138
SCEDC ETAS 81.7% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
SMASH 59.9% 0.602 51.1% 0.471 68.0% 0.611
DSTPP 0.0% 0.840 6.1% 0.935 0.8% 0.988
QTM_SanJac ETAS 88.8% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
SMASH 67.8% 0.290 55.6% 0.385 53.4% 0.584
DSTPP 78.4% 0.110 85.7% 0.240 85.3% 0.136
QTM_SaltonSea ETAS 77.8% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
SMASH 58.6% 0.486 53.6% 0.371 73.7% 0.451
DSTPP 69.6% 0.154 88.8% 0.136 85.6% 0.127
White ETAS 83.4% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
SMASH 61.3% 0.274 84.5% 0.150 67.7% 0.246
DSTPP 0.0% 0.987 30.9% 0.691 32.3% 0.892
ETAS consistently performs best across all datasets and tests. It achieves the highest pass rates and lowest KS statistics, indicating strong calibration and reliability. SMASH shows moderate performance, often outperforming DSTPP but trailing ETAS. Its results vary more across datasets and tests, with occasional strengths (e.g. in White for spatial KS). DSTPP generally performs worse, with lower pass rates and higher KS statistics, especially for the SCEDC and White datasets. However, it achieves relatively good spatial calibration in some cases (e.g., ComCat).
Figure 4: Forecasts from ETAS, SMASH, and DSTPP during the 2010 M7.2 El Mayor-Cucapah earthquake contained in the ComCat dataset. Top: Spatial forecasts for the day following the mainshock. ETAS accurately captures the primary aftershock zone along the Laguna Salada fault system. SMASH produces smoother forecasts with broader spatial spread, while DSTPP concentrates its probability mass north of the mainshock epicenter. Bottom: Cumulative earthquake counts over time, with magnitudes shown as scaled orange circles. Forecast number distributions from each model are plotted with 95% confidence intervals. All models initially underestimate aftershock activity. ETAS and SMASH begin to recover after the first week, whereas DSTPP continues to systematically underpredict event counts throughout the sequence.
5 Discussion and Conclusion
We introduce EarthquakeNPP, a benchmarking platform designed to evaluate Neural Point Process (NPP) models against the state-of-the-art ETAS model for earthquake forecasting. The platform hosts datasets from diverse regions of California, both standard forecasting zones and datasets that incorporate modern detection techniques. We establish two evaluation frameworks tailored to seismology: standard log-likelihood metrics and the generative consistency tests developed by the Collaboratory for the Study of Earthquake Predictability (CSEP), ensuring that successful models can be directly relevant to operational forecasting.
In benchmarking five NPP models against ETAS, we found that none outperformed the baseline, suggesting current NPPs are not yet suitable for operational use. While NPPs often perform well during low-activity background periods, they struggle during active phases following large earthquakes. This may be due to their lack of explicit magnitude dependence — a key feature in ETAS — and their limited memory, as NPPs encode only a fixed-length history. In contrast, ETAS sums over all past events, allowing long-past earthquakes to influence future rates. Future NPPs should incorporate magnitude information and develop mechanisms for longer-term memory.
EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP/tree/main, provides a platform for future NPP developments to be benchmarked against these initial results. The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions. Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
Acknowledgments and Disclosure of Funding
This project is funded by Compass - Centre for Doctoral Training in Computational Statistics and Data Science (EPSRC Grant Ref EP/S023569/1). Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc. This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grant G24AP00059.
Appendix A Earthquake Catalog Data
A.1 Earthquake Catalog Generation
Figure 5: Generating an earthquake catalog involves several key steps: seismic phase picking, magnitude estimation, and the association and location of seismic sources. This process transforms raw waveform data recorded at seismic stations to locations, times, and magnitudes of earthquakes.
Data missingness, referred to in seismology as catalog (in)completeness, is the primary challenge faced with earthquake catalogs. It is an important and unavoidable feature, and is a result of how earthquakes are detected and characterised. Below, we briefly overview the process of generating an earthquake catalog to illustrate the data quality issues. In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecast models.
Seismometers and Seismic Networks. A seismometer is an instrument that detects and records the vibrations caused by seismic waves [75, 71]. It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities. Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., [92] and references therein). High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
From Waveforms to Phase Picking. The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake. Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes [2]. Recent algorithms, such as machine learning classifiers [e.g. 98, 35] and template-matching [e.g. 62], can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake. Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event. Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms [82, 37]. Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist [e.g. 92, and references therein]. Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter [61]. The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers. This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons. This led to introduction of the moment magnitude scale (Mw) [22], which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
$M_{0}=\text{rigidity}\times\text{rupture area}\times\text{slip},$ (7)
where rigidity is a mechanical property of the rock along the fault, rupture area is the area of the fault that slipped, and slip is the distance the fault moved. Mw is determined seismologically via a spectral fitting process to the earthquake waveforms. In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales [e.g. 26, and references therein].
A.2 Earthquake Catalog Completeness
All of the EarthquakeNPP datasets are made publicly available by their respective data centers in raw format. However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing. This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ [e.g., 42, 43].
There are several reasons why an earthquake may not be detected by a seismic network. Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations. Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high [34, 20]. Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g. when phase arrivals of different events overlap at different stations or the amplitudes of small events are swamped by those of large events. Since catalog incompleteness increases for lower magnitude events, typically the task is to find the value $M_{c}$ above which there is approximately $100\%$ detection probability. Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data. Where NPPs have demonstrated an ability to perform well with incomplete data [77], typically a threshold below the completeness biases classical models such as ETAS [67]. Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds [e.g. 40, 77], which we also facilitate in our datasets.
Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law [17], which states that the distribution of earthquake magnitudes follows an exponential probability density function
$f_{GR}(m)=\beta e^{-\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (8)
where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$. Histogram-based approaches, such as the simple Maximum Curvature method [91] as well as many others [e.g. 26, and references therein], identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ [e.g. 65]. During aftershock sequences, $M_{c}(t)$ can be very high [e.g., 1, 19] (See Figure 6a). Thresholding at the maximum value might remove too much data. Instead, modelers either omit particularly incomplete periods during training and testing [32, 21], model the incompleteness itself [25, 89, 55, 18, 19, 45, 20], or accept known biases from disregarding this issue [73]. Spatially, catalogs are less complete farther from the seismic network [42], so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
Figure 6: a) the June 10, 2016 Mw5.2 Borrego Springs earthquake and aftershocks, which occurred on the San Jacinto fault zone and is recorded in the White catalog. An estimate of the magnitude of completeness $M_{c}(t)$ over time using the Maximum Curvature method reveals more incompleteness immediately following the large earthquake. b) magnitude-frequency histograms reveal that truncating the raw White catalog to inside the target region decreases $M_{c}$. Each histogram is fit to the Gutenberg-Richter (GR) law and an estimate of $M_{c}$ for each catalog occurs where the histogram deviates from the (GR) line. c) An estimate of $M_{c}$ for gridded regions of the San Jacinto fault zone, using the raw White catalog.
Appendix B Additional Datasets
Beyond the official EarthquakeNPP datasets, we include 3 further datasets that either provide additional scientific insight or continuity from previous benchmarking works.
Synthetic ETAS Catalogs. We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region. A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al. [59],
$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
where $M$ is the mainshock magnitude. Events below this threshold are removed using mainshocks of Mw 5.2 and above. The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
Deprecated Catalog of Japan. To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al. [5], however with a chronological train-test split and without removing any supposed outlier events. To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
Table 3: Summary of additional datasets, including: magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training events, and number of testing events. The chronological partitioning of training, validation, and testing periods is also detailed. An auxiliary (burn-in) period begins from the "Start" date, followed by the respective starts of the training, validation, and testing periods. All dates are given as 00:$0$ UTC on January $1^{\text{st}}$, unless noted (* refers to 00:$0$ UTC on January $17^{\text{th}}$).
Catalog $\mathbf{M}_{\mathbf{c}}$ Start-Train-Val-Test-End Train Events Test Events
ETAS 2.5 1971-1981-1998-2007-2020 117,550 43,327
ETAS_incomplete 2.5 1971-1981-1998-2007-2020 115,115 42,932
Japan_Deprecated 2.5 1990-1992-2007-2011-2020 22,213 15,368
Figures 7 and 8 report the temporal and spatial log-likelihood scores of all the benchmarked models on additional datasets. On synthetic data generated by the ETAS model the performance of NPPs mirrors the results on the observational data (Figures 2 and 3). The performance of NPPs is more comparable to ETAS in terms of temporal log-likelihood however they cannot capture the distribution of earthquake locations. Change in temporal performance of models between the ETAS and ETAS_incomplete datasets reveal each model’s robustness to the missing data typically present in earthquake catalogs (See section A.2). Auto-STPP and ETAS reduce in performance upon the removal earthquakes during aftershock sequences, whereas DeepSTPP and NSTPP maintain the same performance indicating a robustness to the data missingness.
On the Japan_Deprecated dataset, whilst ETAS remains the best performing model for spatial prediction, for temporal prediction it performs comparably to NSTPP and is even marginally outperformed by DeepSTPP. This performance can be attributed to the data completeness issues of the Japan_Deprecated dataset (see section 1.1), where the test period is missing all earthquakes bellow magnitude 4.0.
Figure 7: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the additional datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 8: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the additional datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Appendix C Computational Efficiency
C.1 Training
Table 4 reports the training times for each model across all datasets. We ran all the NPP models using a HPC node with Nvidia Ampere GPU with 4x Nvidia A100 40GB SXM “Ampere” GPUs and AMD EPYC 7543P 32-Core Processor “Milan” CPU using torch==1.12.0 and cuda==11.3.
Dataset # Training Events ETAS Deep-STPP AutoSTPP NSTPP SMASH DSTPP Poisson
ComCat 79,037 08:59:04 00:15:35 01:34:09 3 days, 05:10:17 2:21:13 20:05:57 <1 second
QTM_SaltonSea 44,042 07:28:28 00:26:46 01:45:34 2 days, 00:26:45 2:38:21 11:12:$0$ <1 second
QTM_SanJac 18,664 00:32:40 00:09:31 00:37:03 1 day, 22:06:33 0:55:34 4:44:46 <1 second
SCEDC_20 128,265 13:42:30 00:38:10 02:54:51 3 days, 02:20:40 4:22:16 1 day, 8:37:05 <1 second
SCEDC_25 43,221 03:09:14 00:09:34 00:56:05 2 days, 16:33:55 1:24:07 10:59:28 <1 second
SCEDC_30 12,426 00:42:25 00:02:44 00:16:01 1 day, 16:39:04 0:24:01 3:10:26 <1 second
White 38,556 03:55:40 00:08:21 01:10:51 2 days, 01:03:57 1:46:17 9:48:47 <1 second
Japan_Deprecated 22,213 06:09:08 00:13:45 01:02:07 2 days, 05:32:03 1:33:06 5:39:32 <1 second
ETAS 117,550 00:33:25 00:15:24 01:10:22 3 days, 03:09:17 1:45:33 1 day, 1:27:44 <1 second
ETAS_incomplete 115,115 00:35:14 00:15:29 01:09:43 3 days, 11:39:51 1:44:34 1 day, 2:28:42 <1 second
[…]
ETAS training scales $\mathcal{O}(n^{2})$ with the total number of events, since for every event a contribution to the intensity function is computed from a summation over all previous events. This scaling, coupled with the lack of parallelization in the current implementation, results in long training times for larger datasets. Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future [76].
Encouragingly, both DeepSTPP and AutoSTPP are significantly faster to train due to GPU acceleration and their use of a sliding window of the most recent $k=20$ events. While exact complexity analyses are not provided in Zhou et al. [97] or Zhou and Yu [96], we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood. AutoSTPP, though requiring automatic integration to compute the likelihood, still scales with $\mathcal{O}(kn)$ because the additional integration cost does not affect the overall scaling.
NSTPP, on the other hand, incurs significant training costs, rendering it impractical for real-time forecasting. Unlike the sliding window mechanism used in DeepSTPP and AutoSTPP, NSTPP partitions the event sequence into fixed time intervals, leading to sequences that are much longer than the $k=20$ events used by the other models (as shown in Figure 11 of Chen et al. [5]). Furthermore, solving an ODE for each event time adds a significant computational burden, even with the use of their faster attentive CNF architecture.
Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP—even faster than ETAS. This efficiency arises from its use of a single-step, normalization-free score-matching objective, which avoids the costly denoising and sampling loops required in diffusion-based training. SMASH directly learns the gradient of the log-density via pseudolikelihood estimation, enabling efficient GPU parallelization and bypassing the need for repeated evaluations over diffusion steps. In contrast, DSTPP simulates a sequential generative process over hundreds of intermediate steps per sample, significantly increasing computation and memory costs.
C.2 Simulation
Real-time earthquake forecasting and CSEP model evaluation require simulating many repeat sequences (at least 10,000 for adequate distributional coverage) over the forecasting horizon. While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation scales more efficiently at $\mathcal{O}(n\log n)$. This improved scaling is due to its equivalent formulation as a Hawkes branching process (see Section 2.2). Both Deep-STPP and AutoSTPP are also based on Hawkes processes, which theoretically allows for fast simulation. However, as these models currently only have an intensity function implementation, simulating events would require a slower thinning procedure [52], limiting their simulation efficiency. NSTPP too has slow simulation, since it must solve an ODE to sample a new event. Due to the slow simulation of DeepSTPP, AutoSTPP and NSTPP, they are not evaluated using the CSEP generative evaluation metrics.
Appendix D Analysis of Likelihood Scores
(a) ComCat
(b) SCEDC
(c) SanJac
(d) SaltonSea
(e) WHITE
Figure 9: Cumulative information gain (IG) plots for the temporal performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
(a) ComCat
(b) SCEDC
(c) SanJac
(d) SaltonSea
(e) WHITE
Figure 10: Cumulative information gain (IG) plots for the spatial performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
Appendix E CSEP Consistency Tests
E.1 Number (Temporal) Test
The number test evaluates the temporal component of the forecast by checking the consistency of the forecasted number of events, $N$ with those observed in the forecast horizon, $N_{\text{obs}}$. Upper and lower quantiles are estimated using the empirical cumulative distribution from the repeat simulations, $F_{N}$,
$\delta_{1}$ $=\mathbb{P}(N\geq N_{\text{obs}})=1-F_{N}(N_{\text{obs}}-1)$ (10)
$\delta_{2}$ $=\mathbb{P}(N\leq N_{\text{obs}})=F_{N}(N_{\text{obs}}).$ (11)
E.2 Pseudo-Likelihood Test
The pseudo-likelihood test evaluates the compatibility of a forecast with an observed catalog using an approximation to the space-time point process likelihood.
The test statistic is based on the pseudo-log-likelihood:
$\hat{L}_{\text{obs}}=\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}_{s}(k_{i})-\bar{N},$ (12)
where $\hat{\lambda}_{s}(k_{i})$ is the approximate rate density in the spatial cell of the $i^{\text{th}}$ event, and $\bar{N}$ is the expected number of events.
Each forecast simulation $j$ provides a test statistic
$\hat{L}_{j}=\sum_{i=1}^{N_{j}}\log\hat{\lambda}_{s}(k_{ij})-\bar{N},$ (13)
which is used to build the empirical cumulative distribution $F_{L}$. The quantile score is then computed as
$\gamma_{L}=\mathbb{P}(\hat{L}_{j}\leq\hat{L}_{\text{obs}})=F_{L}(\hat{L}_{\text{obs}}).$ (14)
E.3 Spatial Test
To evaluate the spatial component of the forecast, a test statistic aggregates the forecasted rates of earthquakes over a regular grid,
$S=\left[\sum_{i=1}^{N}\log\hat{\lambda}(k_{i})\right]N^{-1},$ (15)
where $\hat{\lambda}(k_{i})$ is the approximate rate in the cell $k$ where the $i^{th}$ event is located. Upper and lower quantiles are estimated by comparing the observed statistic
$S_{\text{obs}}=\left[\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}(k_{i})\right]N_{\text{obs}}^{-1},$ (16)
with the empirical cumulative distribution of $S$ using the repeat simulations, $F_{S}$
$\gamma_{s}=\mathbb{P}(S\leq S_{\text{obs}})=F_{S}(S_{\text{obs}}).$ (17)
The grid is constructed from $\{0.1\degree,0.05\degree,0.01\degree\}$ squares for ComCat, SCEDC and {QTM_Salton_Sea, QTM_SanJac, White} respectively.
E.4 Magnitude Test
To evaluate the earthquake magnitude component of the forecast, a test statistic compares the histogram of a forecast’s magnitudes $\Lambda^{(m)}$, against the mean histogram over all forecasts $\bar{\Lambda}^{(m)}$,
$D=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda^{(m)}(k)+1\right]\right)^{2},$ (18)
where $\Lambda^{(m)}(k)$ and $\bar{\Lambda}^{(m)}(k)$ are the counts in the $k^{th}$ bin of the forecast and mean histograms, normalised to have the same total counts as the observed catalog. Upper and lower quantiles are estimated by comparing the observed statistic
$D_{\text{obs}}=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda_{\text{obs}}^{(m)}(k)+1\right]\right)^{2},$ (19)
with the empirical distribution of $D$ using the repeat simulations, $F_{D}$
$\gamma_{m}=\mathbb{P}(D\leq D_{\text{obs}})=F_{D}(D_{\text{obs}}).$ (20)
Histogram bins of size $\delta_{m}=0.1$ are used across all datasets.
(a) Number test
(b) Spatial test
(c) Magnitude test
Figure 11: CSEP consistency tests on the ETAS model for the first day ($01/01/2014$) of the testing period in the SCEDC catalog. A total of 10,000 simulations are generated to compute empirical distributions of the test statistics for each of the three consistency tests: (a) Number test, (b) Spatial test, and (c) Magnitude test. The test fails if the observed statistic falls within the rejection region (red), defined by the $0.05$ and $0.95$ quantiles of the distribution.
E.5 Evaluating Multiple Forecasting Periods
Savran et al. [63] describe how to assess a model’s performance across the multiple days in the testing period (Figure 12). By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator [15]. Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
Figure 12: Daily number forecasts from SMASH on the ComCat dataset. (a) Forecasted daily distributions for the number of earthquakes, with green lines indicating days where the observed count falls within the 95% forecast interval, and red lines where the forecast fails. Observed values are marked with dot sizes proportional to the number of earthquakes. (b) Quantile scores from the number test for each day, with red markers indicating failed forecasts. Marker size reflects the number of earthquakes observed on that day. (c) Temporal evolution of observed earthquakes during the testing period, with event magnitudes represented by marker size. (d) Histogram of quantile scores from the number test. Under ideal calibration, scores should follow a uniform distribution. Red bars indicate failed forecasts, and the Kolmogorov–Smirnov (KS) statistic quantifies deviation from uniformity.
Appendix F Further Dataset Figures
F.1 ComCat
Figure 13: Times and magnitudes of events in the ComCat dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 14: Locations of events in the ComCat dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.2 SCEDC
Figure 15: Times and magnitudes of events in the SCEDC dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 16: Locations of events in the SCEDC dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.3 White
Figure 17: Times and magnitudes of events in the White dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 18: Locations of events in the White dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.4 QTM_SanJac
Figure 19: Times and magnitudes of events in the QTM_SanJac dataset. The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
F.5 QTM_SaltonSea
Figure 20: Times and magnitudes of events in the QTM_SaltonSea dataset. The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 21: Locations of events in the QTM_SanJac and QTM_SaltonSea datasets, labeled by their partition into auxiliary, training, validation and testing periods.
Appendix G Error Distributions & Next-event metrics
Figure 22: The distribution of errors ($Y_{\text{obs}}-Y_{\text{pred}}$) for the Normal(0,1)(0,1), Exponential(1)(1), and Pareto(2)(2) distributions. Maximum likelihood estimation is used to fit Normal and Laplace distributions to each error histogram. Normal errors (Normal $\times$ Normal) are best approximated by the Root Mean Square Error (RMSE), while Laplacian errors (Exponential $\times$ Exponential) are best approximated by the Mean Absolute Error (MAE). However, neither RMSE nor MAE effectively capture the errors for the heavy-tailed Pareto distribution.